# Offline and Online Use of Interval and Set-Based Approaches for Control and State Estimation: A Selection of Methodological Approaches and Their Application

### v3

@@ 1.1. Use of Interval Methods to Deal with Bounded Uncertainty
In its widest sense, interval analysis provides an approach to perform computations on finitely large sets in such a way that the determined solutions contain all the solutions of the mathematically stated problem with certainty [65, 39]. In such a way, it is not only possible to account for numerical errors due to the limited precision of the numbers representable in floating point arithmetic on a CPU, but also for discretization and method errors. The latter arise inevitably when solving, for example, ordinary differential equations with the help of temporal series expansions truncated after a finite order [70, 71, 43, 63, 15]. In addition to these effects, which are small (at least) for a single evaluation step, but can cause severe damage if neglected, the same techniques also allow for computing with wider sets that result from imperfectly known system parameters and initial conditions in (technical) applications. This lack of knowledge is commonly referred to as epistemic uncertainty [51]. As such, interval analysis provides a mathematically sound way to handle epistemic uncertainty that can furthermore also result from other kinds of a lack of knowledge of the actual system behavior, for example, friction phenomena in rigid body-body dynamics or fluidic applications.
@@ 1.3. Challenges of the Use of Set-Based Methods
    $\mathbf{A}^{2}=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}\quad\text{and}\quad\mathbf{A}^{4}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1}{\sqrt{2}}}\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$ (8)
@@ 1.4. Selected Countermeasures Against Overestimation
holds for all elements $i\in\{1,\ldots,n\}$ of the two vectors with $\mathbf{x}^{\langle{1}\rangle}\left({0}\right)$ and $\mathbf{x}^{\langle{2}\rangle}\left({0}\right)$ being the initial conditions which satisfy again the inequalities
@@
Note that couplings between the vectors $\mathbf{v}(t)$ and $\mathbf{w}(t)$ disappear in both summations for $j\neq i$ included in (16) if the system is positive, i.e., $v_{i}(t)\geq 0$ holds for all $t\geq 0$, see for example [74, 25]. There, this property is exploited for the design of a Luenberger-type interval observer, which allows for guaranteed state estimation in the case of bounded measurement uncertainty.
@@ 2.1. Predictor–Corrector Framework for the Verified Identification of Parameters of Dynamic Systems
    • The use of contractors and set inversion approaches to determine an inverse mapping from the measured outputs to the internal system states and parameters [37]. Examples are forward–backward contractors [39] initialized with the predicted state intervals that are intersected with the associated measurement outputs or interval Newton techniques in the case of nonlinear output equations or for output equations involving more than one state variable. Additionally, also observer-based formulation for solving the same task were recently published in [37], where zonotopes are employed as the underlying uncertainty representation.
@@ 3.2. Fundamental Sliding Mode Control Laws
Both the treatment of strict inequality constraints representing the limitations of the admissible operating conditions such as the temperature limit above and bounded interval uncertainty of internal system parameter can be combined with classical first- and second-order sliding mode techniques [26, 31]. These classical approaches have the drawback that they suppose perfectly known parameters and can therefore either not ensure the guaranteed satisfaction of state constraints. Moreover, they may be inefficient in the use of the available ranges of actuator signals [75].
@@ 3.2.1. First-Order Sliding Mode Control
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi}_{1}(t)=x_{1}(t)-x_{1,\mathrm{d}}(t)$ (21)
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi}_{1}^{(r)}(t)=x_{1}^{(r)}(t)-x_{1,\mathrm{d}}^{(r)}(t)~{},~{}~{}r\in\{0,1,\ldots,n\}\enspace,$ (22)
to define the sliding surface [^4] In this context, sliding means that the system dynamics obey a differential equation of lower order smaller than $n$ as soon as the state is bound to the surface $s=0$, cf. [106].
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s(t)=\sum\limits_{r=0}^{n-1}\alpha_{r}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi}_{1}^{(r)}(t)}$ (23)
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V^{\langle{\mathrm{I}}\rangle}(t)=\frac{1}{2}s^{2}(t)}\enspace,$ (24)
@@
$\dot{V}^{\langle{\mathrm{I}}\rangle}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(t)}={s(t)\cdot\dot{s}(t)=\left({\sum\limits_{r=0}^{n-1}\alpha_{r}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi}_{1}^{(r)}(t)}\right)\cdot\left({\sum\limits_{r=0}^{n-1}\alpha_{r}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi}_{1}^{(r+1)}(t)}\right)}<0\enspace,$ (25)
@@
$\left({\sum\limits_{r=0}^{n-1}\alpha_{r}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi}_{1}^{(r)}}\right)\cdot\left({\sum\limits_{r=0}^{n-1}\alpha_{r}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi}_{1}^{(r+1)}}\right)<-\eta\cdot\left|s\right|=-\eta\cdot\left({\sum\limits_{r=0}^{n-1}\alpha_{r}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi}_{1}^{(r)}}\right)\cdot\mathrm{sign}\left({s}\right)$ (26)
@@
The time-independent coefficients $\alpha_{r}$ have a major influence on the control amplitudes if non-modeled errors and disturbances act on the system dynamics and if the error signals ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi}_{1}^{(r)}$ are corrupted by non-negligible measurement noise or state reconstruction errors, preventing the perfect tracking of the desired trajectory $x_{1,\mathrm{d}}(t)$ and making it impossible that $s\equiv 0$ holds perfectly after the end of the reaching phase.
@@
$\sum\limits_{r=0}^{n-2}\alpha_{r}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi}_{1}^{(r+1)}+u-x_{1,\mathrm{d}}^{(n)}+\eta\cdot\mathrm{sign}\left({s}\right)=-\beta\cdot\mathrm{sign}\left({s}\right)\quad\text{with {an arbitrary}}\quad\beta>0\enspace.$ (27)
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u}=u^{\langle{\mathrm{I}}\rangle}=x_{1,\mathrm{d}}^{(n)}-\sum\limits_{r=0}^{n-2}\alpha_{r}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi}_{1}^{(r+1)}-\tilde{\eta}\cdot\mathrm{sign}\left({s}\right)\enspace.$ (28)
@@ 3.2.2. Second-Order Sliding Mode Control
$\gamma_{1}\dot{s}+\gamma_{0}s=\sum\limits_{r=-1}^{n-1}\alpha_{r}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi}_{1}^{(r)}\quad\text{with}\quad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi}_{1}^{(-1)}(t):=\int_{0}^{t}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi}_{1}(\tau)\mathrm{d}\tau\,.$ (29)
For the sake of asymptotic stability, the coefficients $\gamma_{0}$ and $\gamma_{1}$ need to be strictly positive. For $\alpha_{-1}=0$, this sliding surface has a proportional and differentiating characteristic, while $\alpha_{-1}\neq 0$ means that the integral ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi}_{1}^{(-1)}$ of the output error is additionally fed back in the definition of the sliding variable $s$.
@@
$\dot{V}^{\langle{\mathrm{II}}\rangle}=s\cdot\dot{s}+\dot{s}\cdot\left({-\frac{\lambda\gamma_{0}}{\gamma_{1}}\dot{s}+\frac{\lambda}{\gamma_{1}}\sum\limits_{r=0}^{n}\alpha_{r-1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi}_{1}^{(r)}}\right)<0\enspace,$ (31)
@@
$u=u^{\langle{\mathrm{II}}\rangle}=x_{1,\mathrm{d}}^{(n)}+\frac{1}{\alpha_{n-1}}\cdot\Biggl{(}\gamma_{0}\dot{s}-s-\sum\limits_{r=0}^{n-1}\alpha_{r-1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi}_{1}^{(r)}-\mathrm{sign}\left({\dot{s}}\right)\cdot\left({\tilde{\eta}_{1}+\tilde{\eta}_{2}\cdot\left|{s}\right|}\right)\Biggr{)}$ (33)
@@ 3.3.1. One-Sided State Constraints
$s\cdot\underbrace{\left({\sum\limits_{r=0}^{n-2}\alpha_{r}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi}_{1}^{(r+1)}+u-x_{1,\mathrm{d}}^{(n)}+\eta\cdot\mathrm{sign}\left({s}\right)+\frac{1}{s}\cdot\dot{V}^{\langle{\mathrm{A}}\rangle}}\right)}_{-\beta\cdot\mathrm{sign}\left({s}\right)}<0$ (38)
@@ 3.4. Interval Extensions to Handle Bounded Parameter Uncertainty and State Estimation Errors
${0\not\in\left[{b}\right]\left({\left[{\mathbf{x}}\right],\left[{\mathbf{p}}\right]}\right)}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},\quad\forall{{\mathbf{x}}\in\left[{{\mathbf{x}}}\right]}\quad\text{and}\quad\forall{\mathbf{p}}\in\left[{{\mathbf{p}}}\right]}$ (46)
@@
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi}_{1}^{(r)}(t)\in\left[{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi}_{1}^{(r)}}\right](t)=\left[{x_{1}^{(r)}}\right](t)-x_{1,\mathrm{d}}^{(r)}(t)}$ (47)
@@
$\left[{v^{\langle{\mathrm{I}}\rangle}}\right]={\tfrac{-\left[{a}\right]\left({\left[{{\mathbf{x}}}\right],\left[{{\mathbf{p}}}\right]}\right)+x_{1,\mathrm{d}}^{(n)}-\sum\limits_{r=0}^{n-2}\alpha_{r}\cdot\left[{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi}_{1}^{(r+1)}}\right]-\tilde{\eta}\cdot\mathrm{sign}\left({\left[{s}\right]}\right)}{\left[{b}\right]\left({\left[{\mathbf{x}}\right],\left[{\mathbf{p}}\right]}\right)}}\enspace,$ (48)

### v4

@@ 1.1. Use of Interval Methods to Deal with Bounded Uncertainty
In its widest sense, interval analysis provides an approach to perform computations on finitely large sets in such a way that the determined solutions contain all the solutions of the mathematically stated problem with certainty [65, 39]. In such a way, it is not only possible to account for numerical errors due to the limited precision of the numbers representable in floating point arithmetic on a CPU, but also for discretization and method errors. The latter arise inevitably when solving, for example, ordinary differential equations with the help of temporal series expansions truncated after a finite order [70, 71, 43, 63, 15]. In addition to these effects, which are small (at least) for a single evaluation step, but can cause severe damage if neglected, the same techniques also allow for computing with wider sets that result from imperfectly known system parameters and initial conditions in (technical) applications. This lack of knowledge is commonly referred to as epistemic uncertainty [51]. As such, interval analysis provides a mathematically sound way to handle epistemic uncertainty that can furthermore also result from other kinds of a lack of knowledge of the actual system behavior, for example, friction phenomena in dynamics of rigid bodies or fluidic applications.
@@ 1.3. Challenges of the Use of Set-Based Methods
    $\mathbf{A}^{2}=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}\quad\text{and}\quad\mathbf{A}^{4}={\frac{1}{\sqrt{2}}}\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$ (8)
@@ 1.4. Selected Countermeasures Against Overestimation
holds at each $t\geq t_{0}$ for all elements $i\in\{1,\ldots,n\}$ of the two vectors with $\mathbf{x}^{\langle{1}\rangle}\left({t_{0}}\right)$ and $\mathbf{x}^{\langle{2}\rangle}\left({t_{0}}\right)$ being the initial conditions which satisfy again the inequalities
@@
Note that couplings between the vectors $\mathbf{v}(t)$ and $\mathbf{w}(t)$ disappear in both summations for $j\neq i$ included in (16) if the system is positive, i.e., $v_{i}(t)\geq 0$ holds for all $t\geq t_{0}$, see for example [74, 25]. There, this property is exploited for the design of a Luenberger-type interval observer, which allows for guaranteed state estimation in the case of bounded measurement uncertainty.
@@ 2.1. Predictor–Corrector Framework for the Verified Identification of Parameters of Dynamic Systems
    • The use of contractors and set inversion approaches to determine an inverse mapping from the measured outputs to the internal system states and parameters [37]. Examples are forward–backward contractors [39] initialized with the predicted state intervals that are intersected with the associated measurement outputs or interval Newton techniques in the case of nonlinear output equations or for output equations involving more than one state variable. Additionally, also observer-based formulations for solving the same task were recently published in [37], where zonotopes are employed as the underlying uncertainty representation.
@@ 3.2. Fundamental Sliding Mode Control Laws
Both the treatment of strict inequality constraints representing the limitations of the admissible operating conditions such as the temperature limit above and bounded interval uncertainty of internal system parameters can be combined with classical first- and second-order sliding mode techniques [26, 31]. These classical approaches have the drawback that they suppose perfectly known parameters and can therefore typically not ensure the guaranteed satisfaction of state constraints. Moreover, they may be inefficient in the use of the available ranges of actuator signals [75].
@@ 3.2.1. First-Order Sliding Mode Control
${\xi}_{1}(t)=x_{1}(t)-x_{1,\mathrm{d}}(t)$ (21)
@@
${\xi}_{1}^{(r)}(t)=x_{1}^{(r)}(t)-x_{1,\mathrm{d}}^{(r)}(t)~{},~{}~{}r\in\{0,1,\ldots,n\}\enspace,$ (22)
to define the sliding surface [^4] In this context, sliding means that the system dynamics obey a differential equation of order smaller than $n$ as soon as the state is bound to the surface $s=0$, cf. [106].
${s(t)=\sum\limits_{r=0}^{n-1}\alpha_{r}{\xi}_{1}^{(r)}(t)}$ (23)
@@
${V^{\langle{\mathrm{I}}\rangle}(t)=\frac{1}{2}s^{2}(t)}\enspace,$ (24)
@@
$\dot{V}^{\langle{\mathrm{I}}\rangle}{(t)}={s(t)\cdot\dot{s}(t)=\left({\sum\limits_{r=0}^{n-1}\alpha_{r}{\xi}_{1}^{(r)}(t)}\right)\cdot\left({\sum\limits_{r=0}^{n-1}\alpha_{r}{\xi}_{1}^{(r+1)}(t)}\right)}<0\enspace,$ (25)
@@
$\left({\sum\limits_{r=0}^{n-1}\alpha_{r}{\xi}_{1}^{(r)}}\right)\cdot\left({\sum\limits_{r=0}^{n-1}\alpha_{r}{\xi}_{1}^{(r+1)}}\right)<-\eta\cdot\left|s\right|=-\eta\cdot\left({\sum\limits_{r=0}^{n-1}\alpha_{r}{\xi}_{1}^{(r)}}\right)\cdot\mathrm{sign}\left({s}\right)$ (26)
@@
The time-independent coefficients $\alpha_{r}$ have a major influence on the control amplitudes if non-modeled errors and disturbances act on the system dynamics and if the error signals ${\xi}_{1}^{(r)}$ are corrupted by non-negligible measurement noise or state reconstruction errors, preventing the perfect tracking of the desired trajectory $x_{1,\mathrm{d}}(t)$ and making it impossible that $s\equiv 0$ holds perfectly after the end of the reaching phase.
@@
$\sum\limits_{r=0}^{n-2}\alpha_{r}{\xi}_{1}^{(r+1)}+u-x_{1,\mathrm{d}}^{(n)}+\eta\cdot\mathrm{sign}\left({s}\right)=-\beta\cdot\mathrm{sign}\left({s}\right)\quad\text{with {an arbitrary}}\quad\beta>0\enspace.$ (27)
@@
${u}=u^{\langle{\mathrm{I}}\rangle}=x_{1,\mathrm{d}}^{(n)}-\sum\limits_{r=0}^{n-2}\alpha_{r}{\xi}_{1}^{(r+1)}-\tilde{\eta}\cdot\mathrm{sign}\left({s}\right)\enspace.$ (28)
@@ 3.2.2. Second-Order Sliding Mode Control
$\gamma_{1}\dot{s}+\gamma_{0}s=\sum\limits_{r=-1}^{n-1}\alpha_{r}{\xi}_{1}^{(r)}\quad\text{with}\quad{\xi}_{1}^{(-1)}(t):=\int_{0}^{t}{\xi}_{1}(\tau)\mathrm{d}\tau\,.$ (29)
For the sake of asymptotic stability, the coefficients $\gamma_{0}$ and $\gamma_{1}$ need to be strictly positive. For $\alpha_{-1}=0$, this sliding surface has a proportional and differentiating characteristic, while $\alpha_{-1}\neq 0$ means that the integral ${\xi}_{1}^{(-1)}$ of the output error is additionally fed back in the definition of the sliding variable $s$.
@@
$\dot{V}^{\langle{\mathrm{II}}\rangle}=s\cdot\dot{s}+\dot{s}\cdot\left({-\frac{\lambda\gamma_{0}}{\gamma_{1}}\dot{s}+\frac{\lambda}{\gamma_{1}}\sum\limits_{r=0}^{n}\alpha_{r-1}{\xi}_{1}^{(r)}}\right)<0\enspace,$ (31)
@@
$u=u^{\langle{\mathrm{II}}\rangle}=x_{1,\mathrm{d}}^{(n)}+\frac{1}{\alpha_{n-1}}\cdot\Biggl{(}\gamma_{0}\dot{s}-s-\sum\limits_{r=0}^{n-1}\alpha_{r-1}{\xi}_{1}^{(r)}-\mathrm{sign}\left({\dot{s}}\right)\cdot\left({\tilde{\eta}_{1}+\tilde{\eta}_{2}\cdot\left|{s}\right|}\right)\Biggr{)}$ (33)
@@ 3.3.1. One-Sided State Constraints
$s\cdot\underbrace{\left({\sum\limits_{r=0}^{n-2}\alpha_{r}{\xi}_{1}^{(r+1)}+u-x_{1,\mathrm{d}}^{(n)}+\eta\cdot\mathrm{sign}\left({s}\right)+\frac{1}{s}\cdot\dot{V}^{\langle{\mathrm{A}}\rangle}}\right)}_{-\beta\cdot\mathrm{sign}\left({s}\right)}<0$ (38)
@@ 3.4. Interval Extensions to Handle Bounded Parameter Uncertainty and State Estimation Errors
${0\not\in\left[{b}\right]\left({\left[{\mathbf{x}}\right],\left[{\mathbf{p}}\right]}\right)}{,\quad\forall{{\mathbf{x}}\in\left[{{\mathbf{x}}}\right]}\quad\text{and}\quad\forall{\mathbf{p}}\in\left[{{\mathbf{p}}}\right]}$ (46)
@@
${{\xi}_{1}^{(r)}(t)\in\left[{{\xi}_{1}^{(r)}}\right](t)=\left[{x_{1}^{(r)}}\right](t)-x_{1,\mathrm{d}}^{(r)}(t)}$ (47)
@@
$\left[{v^{\langle{\mathrm{I}}\rangle}}\right]={\tfrac{-\left[{a}\right]\left({\left[{{\mathbf{x}}}\right],\left[{{\mathbf{p}}}\right]}\right)+x_{1,\mathrm{d}}^{(n)}-\sum\limits_{r=0}^{n-2}\alpha_{r}\cdot\left[{{\xi}_{1}^{(r+1)}}\right]-\tilde{\eta}\cdot\mathrm{sign}\left({\left[{s}\right]}\right)}{\left[{b}\right]\left({\left[{\mathbf{x}}\right],\left[{\mathbf{p}}\right]}\right)}}\enspace,$ (48)